Inverse Laplace Transforms

Find L^-1 {F(s)} when F(s) is given by


1. s+5/(s+1)(s-3)


1
Expert's answer
2022-05-25T17:10:46-0400

1.


"\dfrac{s+5}{(s+1)(s-3)}=\dfrac{A}{s+1}+\dfrac{B}{s-3}"



"=\dfrac{As-3A+Bs+B}{(s+1)(s-3)}"


"A+B=1""-3A+B=5"




"A=-1, B=2"




"L^{-1}(\dfrac{s+5}{(s+1)(s-3)})=-L^{-1}(\dfrac{1}{s+1})"




"+2L^{-1}(\dfrac{1}{s-3})=-e^{-t}+2e^{3t}"